Q6: ažu au äx2 + 0. 0 < x <a, 0 < y < b 022 u(0. y) = 1. ut, y) = 1 u(x, 0) = 0, u(x, TT) = |

solve this
Answer #1

2²ų ahu + ocycb } we get =) = We. 86 alu Ocula 2x2 dyz U(0.9) = 1 i U(1.7) = 1 u(1,0) = U(X, 1) = 1 By Assuming variable sepaYo=0 M C cosh (o) + D sinh (o) = o h. I a C. 1 D.00 Te=0 [:c=o And YM) = 1 I sinh (H&) = 1 > D sinh (Hd) = 1 Boundary conditiand Y) = ie Y ) = e cosh (2nn) to sinh (20x) 1. sinh (2nx] [:c=0,D=1+0] therefore the solution of this Laplace Equation is uc